Solve the given initial-value problem. The DE is homogeneous. xv2 dx = y3 – x3, 7(1) = 4 y=x(64 – In(x))(+)
Answer #1

Given D.E. isi ny dy - y3 w3 dn y 3 x3 dy dn nya Clearly put above DE is Homo. y=va - dy - xt ndy dan We get, SAKA) up k te u

which is in variable separable form. Integrating above eq we get . fredx+ ft dn ac M + enim) - C put v= 4x +3 en (n) = 36. 21by using definition of solution of homogeneous differential eq i was solved this problem